One pump can empty a tank three times as another. working together they can complete a job in four hours, find the rate of each pump
Answer by stanbon(75887)

One pump can empty a tank three times as another. working together they can complete a job in four hours, find the rate of each pump
---
One pump rate: 1/x job/hr
Other pump rate: 1/(3x) job/hr
-----
Together rate: 1/4 job/hr
================================
Equation:
rate + rate = together rate
1/x + 1/(3x) = 1/4
---
Multiply thru by 12x to get:
12 + 4 = 3x
3x = 16
-----
One pump DATA:
x = 16/3 hr
Pump rate = (3/16) job/hr
-----
Other pump DATA:
time = 3x = 16 hr
Pump rate = 1/16 job/hr
=============================
Cheers,
Stan H.
============